OpenStudy (anonymous):

Tom receives allowance every week. He spends three-fourths of his allowance on food. Last week, Tom spent nine dollars on food. What is Tom’s weekly allowance? Write the verbal model, define the variable, write the equation and find the solution. Show your work. *I am stuck on how to make it into a equation and its suppose to be a one step equation*

OpenStudy (jackellyn):

3/4x=9 3/4 Because he spends 3/4 of his allowance so x is his allowance they tell you he spent 9 dollars. 4/3(9)=x 12=x so his allowance is 12 dollars you can check 12*(3/4)=9 which is what he spent last week. understand @poppur?